# Distillation and Refinement of Reasoning in Small Language Models for Document Re-ranking

### v2

@@ Abstract.
journalyear: 2025copyright: rightsretainedccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Search engines and retrieval-augmented generation systems increasingly face queries that require complex reasoning and multi-step synthesis and analysis. They demand a deeper understanding of the query and the documents to identify the connections between them. For example, finding documentation for a coding error requires understanding program logic and syntax, and identifying economic case studies that share underlying theoretical principles demands sophisticated domain knowledge and analytical reasoning (Su et al., 2024). Traditional approaches to training ranking models for such complex tasks often rely on expensive human annotations to provide relevance judgments and explanations. In contrast, we present a framework that automatically generates its own training signal by leveraging existing question-answer pairs on the Web.
Although neural ranking models have made significant progress in recent years (Nogueira and Cho, 2019; Nogueira et al., 2019; Devlin et al., 2018; Samarinas et al., 2021; Killingback et al., 2025) and led to substantial performance gains on standard benchmarks such as MS MARCO (Nguyen et al., 2016) and the TREC Deep Learning (DL) Track (Craswell et al., 2021, 2020), we observe that they often struggle with reasoning-intensive queries that demand deeper understanding and explicit justification of relevance decisions. For instance, state-of-the-art dense retrievers that achieve strong performance on TREC DL show significant degradation on reasoning-intensive queries, with the best models achieving only about 18% nDCG@10 on the BRIGHT benchmark (Su et al., 2024)–a recent benchmark designed for reasoning-intensive ranking tasks. We argue that ranking models must engage in deliberate reasoning to bridge the gap between query intent and document content.
@@ 2.1. Model Architecture
Second-stage Re-ranking: Various learning-to-rank models, from traditional feature-based (Liu, 2010; Burges, 2010) to transformer-based cross-encoder models (Nogueira and Cho, 2019; Nogueira et al., 2019; Gao et al., 2021), have been used for reranking. We aim at training a reranking model for effective reasoning-intensive tasks. To do so, we train a language model that takes a query-document pair at a time and generates some reasoning to analyze and describe whether and how the provided document is relevant. These reasoning steps are then followed by a discrete relevance label as the final generation token. This relevance label is either 0 (i.e., non-relevant), 1 (partially relevant), or 2 (highly relevant). This stage is crucial for complex reasoning tasks, as it allows deeper analysis of document content in relation to the query intent. Since the scores produced by our reranker are discrete, many documents are assigned the same relevance score, and we cannot distinguish them in ranking. Therefore, we employ a hybrid scoring strategy that combines the generated discrete reranking score with the retrieval score produced in the first-stage retrieval. In fact,
@@ 2.2. Model Optimization
1. Synthetic Data Generation: High-quality training data is crucial for developing models that can handle diverse reasoning patterns. However, obtaining human annotations for reasoning-intensive ranking is expensive and time-consuming. We address this challenge through an automated data generation process that leverages existing question-answer pairs from social websites like StackExchange. Our data generation pipeline, summarized in Algorithm 1 and Figure 1, starts with a seed set of query-answer pairs $\mathcal{C}$. In our experiments, we sampled 20K pairs round robin from 186 different communities on StackExchange. For each answer, we extract linked documents (hyperlinks) that potentially contain supporting evidence, establishing an initial set of query-document pairs. To increase diversity and to source potential negative documents, we use a teacher LLM to generate related queries and retrieve additional documents through web search using the Brave Search API. The teacher model is then instructed to generate an explanation and a discrete relevance label for each query-document pair, creating a distillation dataset. This approach naturally captures diverse reasoning patterns since the teacher model must explain how different types of evidence support or fail to support answers across technical domains - from code analysis to scientific explanations. The explanations demonstrate different forms of reasoning like logical deduction, causal analysis, and domain-specific technical reasoning. The next two phases are summarized in Algorithm 2.
@@
15:   $D_{\text{synth}}\leftarrow D_{\text{synth}}\cup\{(q,d,e,l)\}$
@@ 3. Experiments
Our evaluation uses the BRIGHT benchmark (Su et al., 2024), which spans diverse domains requiring complex reasoning capabilities. BRIGHT includes seven datasets from StackExchange communities (Biology, Earth Science, Economics, Psychology, Robotics, Stack Overflow, and Sustainable Living), each containing 100-200 expert-validated query-document pairs where relevance is determined by citations in accepted answers. The remaining 5 datasets focus on coding and mathematical reasoning: Pony (syntax documentation pairs), LeetCode (algorithmic problems), TheoremQA (theorem-based questions), AoPS (competition math problems), and Theorem Retrieval (problems paired with ProofWiki statements). In total, BRIGHT contains 1,384 queries with 6.37 positive documents per query on average. The queries are typically long-form questions requiring multi-step reasoning, while positive documents provide critical concepts, theories, or techniques needed to address the queries rather than direct answers.
@@
The base LLM for InteRank is Llama 3.2 3B, while Llama 3.3 70B is our teacher model (Grattafiori et al., 2024). We use QLoRA (Dettmers et al., 2024) for parameter-efficient fine-tuning, with a 4-bit quantization of the base model and trainable rank-64 adapters. Due to resource constraints, we limit the context length to 4K tokens. Training is performed on a single A100 GPU with an effective batch size of 16 (batch size 1 with 16 gradient accumulation steps) using the AdamW optimizer with learning rate 2e-4. For the sampling of outputs in the RL stage, we use temperature 1.0 for nucleus sampling, reward threshold $\tau=0.85$, and reward scaling power $m=3$. We perform two epochs of RL training. For the reward model, we use a pretrained Llama 3.1 8B model (Liu et al., 2024) that has demonstrated strong performance on RewardBench (Lambert et al., 2024). We found that this model has very high agreement with larger open-weight and commercial LLMs in relative comparison of explanation outputs, making it suitable for our training process.

### v3

@@ Abstract.
reasoning intensive retrieval; reinforcement learning
journalyear: 2025copyright: ccconference: Proceedings of the 2025 International ACM SIGIR Conference on Innovative Concepts and Theories in Information Retrieval (ICTIR); July 18, 2025; Padua, Italybooktitle: Proceedings of the 2025 International ACM SIGIR Conference on Innovative Concepts and Theories in Information Retrieval (ICTIR) (ICTIR ’25), July 18, 2025, Padua, Italydoi: 10.1145/3731120.3744613isbn: 979-8-4007-1861-8/2025/07ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Search engines and retrieval-augmented generation systems increasingly face queries that require complex reasoning and multi-step synthesis and analysis. They demand a deeper understanding of the query and the documents to identify the connections between them. For example, finding documentation for a coding error requires understanding program logic and syntax, and identifying economic case studies that share underlying theoretical principles demands sophisticated domain knowledge and analytical reasoning (SU et al., 2025). Traditional approaches to training ranking models for such complex tasks often rely on expensive human annotations to provide relevance judgments and explanations. In contrast, we present a framework that automatically generates its own training signal by leveraging existing question-answer pairs on the Web.
Although neural ranking models have made significant progress in recent years (Nogueira and Cho, 2019; Nogueira et al., 2019; Devlin et al., 2018; Samarinas et al., 2021; Killingback et al., 2025) and led to substantial performance gains on standard benchmarks such as MS MARCO (Nguyen et al., 2016) and the TREC Deep Learning (DL) Track (Craswell et al., 2021, 2020), we observe that they often struggle with reasoning-intensive queries that demand deeper understanding and explicit justification of relevance decisions. For instance, state-of-the-art dense retrievers that achieve strong performance on TREC DL show significant degradation on reasoning-intensive queries, with the best models achieving only about 18% nDCG@10 on the BRIGHT benchmark (SU et al., 2025)–a recent benchmark designed for reasoning-intensive ranking tasks. We argue that ranking models must engage in deliberate reasoning to bridge the gap between query intent and document content.
@@ 2.1. Model Architecture
Second-stage Re-ranking: Various learning-to-rank models, from traditional feature-based (Liu, 2010; Burges, 2010) to transformer-based cross-encoder models (Nogueira and Cho, 2019; Nogueira et al., 2019; Gao et al., 2021), have been used for reranking. We aim at training a reranking model for effective reasoning-intensive tasks. To do so, we train a language model that takes a query-document pair at a time and generates some reasoning to analyze and describe whether and how the provided document is relevant. These reasoning steps are then followed by a discrete relevance label as the final generation token. This relevance label is either 0 (i.e., non-relevant), 1 (partially relevant), or 2 (highly relevant). This stage is crucial for complex reasoning tasks, as it allows deeper analysis of document content in relation to the query intent. Since the scores produced by our reranker are discrete, many documents are assigned the same relevance score, and we cannot distinguish them in ranking. Therefore, we employ a hybrid scoring strategy that combines the generated discrete reranking score with the (dot product) retrieval score produced in the first-stage retrieval. In fact,
@@ 2.2. Model Optimization
1. Synthetic Data Generation: High-quality training data is crucial for developing models that can handle diverse reasoning patterns. However, obtaining human annotations for reasoning-intensive ranking is expensive and time-consuming. We address this challenge through an automated data generation process that leverages existing question-answer pairs from social websites like StackExchange. Our data generation pipeline, summarized in Algorithm 1 and Figure 1, starts with a seed set of query-answer pairs $\mathcal{C}$. In our experiments, we sampled 20K pairs round robin from 186 different communities on StackExchange. To prevent data contamination, we ensure that any examples appearing in the BRIGHT benchmark from StackExchange are excluded from our training data. For each answer, we extract linked documents (hyperlinks) that potentially contain supporting evidence, establishing an initial set of query-document pairs. To increase diversity and to source potential negative documents, we use a teacher LLM to generate related queries and retrieve additional documents through web search using the Brave Search API. The teacher model is then instructed to generate an explanation and a discrete relevance label for each query-document pair, creating a distillation dataset. This approach naturally captures diverse reasoning patterns since the teacher model must explain how different types of evidence support or fail to support answers across technical domains - from code analysis to scientific explanations. The explanations demonstrate different forms of reasoning like logical deduction, causal analysis, and domain-specific technical reasoning. The next two phases are summarized in Algorithm 2.
@@
15:   $D_{\text{synth}}\leftarrow D_{\text{synth}}\cup\{(q^{\prime},d,e,l)\}$
@@ 3. Experiments
Our evaluation uses the BRIGHT benchmark (SU et al., 2025), which spans diverse domains requiring complex reasoning capabilities. BRIGHT includes seven datasets from StackExchange communities (Biology, Earth Science, Economics, Psychology, Robotics, Stack Overflow, and Sustainable Living), each containing 100-200 expert-validated query-document pairs where relevance is determined by citations in accepted answers. The remaining 5 datasets focus on coding and mathematical reasoning: Pony (syntax documentation pairs), LeetCode (algorithmic problems), TheoremQA (theorem-based questions), AoPS (competition math problems), and Theorem Retrieval (problems paired with ProofWiki statements). In total, BRIGHT contains 1,384 queries with 6.37 positive documents per query on average. The queries are typically long-form questions requiring multi-step reasoning, while positive documents provide critical concepts, theories, or techniques needed to address the queries rather than direct answers.
@@
The base LLM for InteRank is Llama 3.2 3B, while Llama 3.3 70B is our teacher model (Grattafiori et al., 2024). We use QLoRA (Dettmers et al., 2024) for parameter-efficient fine-tuning, with a 4-bit quantization of the base model and trainable rank-64 adapters. Due to resource constraints, we limit the context length to 4K tokens. Training is performed on a single A100 GPU with an effective batch size of 16 (batch size 1 with 16 gradient accumulation steps) using the AdamW optimizer with learning rate 2e-4. For the sampling of outputs in the RL stage, we use temperature 1.0 for nucleus sampling, reward threshold $\tau=0.85$, and reward scaling power $m=3$. We perform two epochs of RL training. For the reward model, we use a pretrained Llama 3.1 8B model (Liu et al., 2024) [^2] https://hf.co/Skywork/Skywork-Reward-Llama-3.1-8B-v0.2 that has demonstrated strong performance on RewardBench (Lambert et al., 2024). We found that this model has very high agreement with larger open-weight and commercial LLMs in relative comparison of explanation outputs, making it suitable for our training process.